OpenStudy (anonymous):

rationalize the denominator of the square root of 9x over 2

OpenStudy (anonymous):

\[(3\sqrt{x})/\sqrt{2}\]

OpenStudy (anonymous):

doesn't really rationalize the denominator though, does it? \(\sqrt2\) is not a rational number

OpenStudy (anonymous):

((3(2)^1/2(x)^1/2))/2

OpenStudy (anonymous):

\[\sqrt{\frac{9x}{2}}=\frac{\sqrt{9x}}{\sqrt2}=\frac{3\sqrt{x}}{\sqrt{2}}\] \[=\frac{3\sqrt{x}}{\sqrt2}\times \frac{\sqrt2}{\sqrt2}=\frac{3\sqrt{2x}}{2}\] now it is rationalized

OpenStudy (anonymous):

got it right this time,was a mistake. too fast